wawa305509:

The acceleration of an object due to gravity is 32 feet per second squared. What is acceleration due to gravity in inches per second squared?

tamyac:

the answer is One foot = 12 inches. 32 feet = x inches x/12=32/1 so x=12*32/1 inches per second^2